There are two masses m1and m2 placed at a distance l apart let the centre of mass of this system is at point named C. If m1 is displaced by l1 towards C and m2 is displaced away from C, find the distance from C where the new centre of mass will be located.

Grade:11

1 Answers

Vikas TU
7 years ago
Initially at centre of mass C,
let x be the distance for m1 and L-x for m2 mass from centre of mass C.
m1x = m2(L-x)...........(1)
After getting displaced the eqn. would form,
m1(L1 + x – h) = m2(L – x + h + L2)...........(2)
solving eqns. (1) and (2) we get,
m1L1 +m1x – m1h = m2L - m2x + m2L2 + m2h
h(m1 + m2) = (m1 + m2)x + m1L1 – m2L – m2L2
h = (m1L1 – m2L)/(m1 +m2)